OpenStudy (anonymous):

Use even and odd properties of the trigonometric functions to find the exact value of the expression.

OpenStudy (anonymous):

csc -( pi / 6 )

OpenStudy (anonymous):

Well, here we go: You started with: \[\csc (-\frac{ \pi }{ 6 })\] Is that right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Now, what is equal to csc (x)? (In terms of another trig function)

OpenStudy (anonymous):

1/sin?

OpenStudy (anonymous):

Okay good, now we get this: \[\csc (-\frac{ \pi }{ 6 }) = \frac{ 1 }{ \sin(-\frac{ \pi }{ 6 }) }\] Correct?

OpenStudy (anonymous):

Yes,

OpenStudy (anonymous):

So what do you get for: \[\sin (-\frac{ \pi }{ 6 })\]

OpenStudy (anonymous):

-.5

OpenStudy (anonymous):

Great, now what is 1/(-1/2)?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

And that's your answer :D

OpenStudy (anonymous):

Okay, thank you for the help.! :-)

OpenStudy (anonymous):

No problem kiddo.